# Network Alignment with Transferable Graph Autoencoders

### v2

@@ 1 Introduction
Network alignment, also known as graph matching, is a classical problem in graph theory, that aims to find node correspondence across different graphs and is vital in a number of high-impact domains (Emmert-Streib et al., 2016). In social networks, for instance, network alignment has been used for user deanonymization (Nilizadeh et al., 2014) and analysis (Ogaard et al., 2013), while in bioinformatics it is a key tool to identify functionalities in protein complexes (Singh et al., 2008), or to identify gene–drug modules (Chen et al., 2018). Graph matching also finds application in computer vision (Conte et al., 2003), sociology (Racz & Sridhar, 2021), or politics (Li et al., 2022), to name a few. Graph matching can be cast as a quadratic assignment problem (QAP), which is in general NP-hard (Koopmans & Beckmann, 1957).
Various approaches have been developed to tackle network alignment and can be divided into two main categories; i) optimization algorithms that attempt to approximate the QAP problem by relaxing the combinatorial constraints, ii) embedding methods that approach the problem by implicitly or explicitly generating powerful node embeddings that facilitate the alignment task. Optimization approaches, as (Anstreicher & Brixius, 2001; Vogelstein et al., 2015) employ quadratic programming relaxations, while (Klau, 2009) and (Peng et al., 2010) utilize semidefinite or Lagrangian-based relaxations respectively. Successive convex approximations were also proposed by (Konar & Sidiropoulos, 2020) to handle the QAP. Challenges associated with these methods include high computational cost, infeasible solutions, or nearly optimal initialization requirements. Embedding methods, on the other hand, overcome these challenges, but they usually produce inferior solutions, due to an inherent trade-off between embedding permutation-equivariance and the ability to capture the structural information of the graph. Typical embedding techniques include spectral and factorization methods (Umeyama, 1988; Feizi et al., 2019; Zhang & Tong, 2016; Kanatsoulis & Sidiropoulos, 2022), structural feature engineering methods (Berlingerio et al., 2013; Heimann et al., 2018), and random walk approaches (Perozzi et al., 2014; Grover & Leskovec, 2016a). Recently (Chen et al., 2020; Karakasis et al., 2021) have proposed joint node embedding and network alignment, to overcome these challenges, but these methods do not scale up as the size of the graph increases.
@@
Graph Neural Networks (GNNs) are powerful architectures that learn graph representations (embeddings). They have shown state-of-the-art performance in several tasks, including biology (Gainza et al., 2020; Strokach et al., 2020; Jiang et al., 2021), quantum chemistry (Gilmer et al., 2017), social networks and recommender systems (Ying et al., 2018; Wu et al., 2020). Recently, (Gao et al., 2021) proposed a GNN approach to match attributed graphs. The method used a joint embedding framework for pairs of graphs and achieved high levels of matching accuracy. However, this method does not scale to large graphs, since training graphs with large sizes is computationally prohibitive.
@@ 2.1 Network Alignment
As mentioned in the introduction, network alignment, is equivalent to the QAP, which has been proven to be NP-hard (Koopmans & Beckmann, 1957).
@@ 3 Graph Neural Networks (GNNs) Upper-Bounds Spectral Methods for Network Alignment
The above property is not satisfied by other spectral methods. GNNs are also stable (Gama et al., 2020), transferable (Ruiz et al., 2020), and have high expressive power (Xu et al., 2019b; Abboud et al., 2021; Kanatsoulis & Ribeiro, 2022).
@@ 3.1 GNNs and Network Alignment
The proof can be found in Appendix C. The assumption that the graph adjacencies have different eigenvalues is not restrictive. Real nonisomorphic graphs have different eigenvalues with very high probability (Haemers & Spence, 2004). Theorem 3.2 compares the network alignment power of a GNN with that of a spectral algorithm (Umeyama, 1988), that uses the absolute values of graph adjacency eigenvectors to match two different graphs. According to Theorem 3.2 there always exists a GNN that can perform at least as well as the spectral approach. The proof studies a GNN with white random input and measures the variance of the filter output. Then it shows that GNN layers are able to compute the absolute values of the graph adjacency eigenvectors when the adjacency has non-repeated eigenvalues. As a result there always exists a single layer GNN that outputs the same node features as the ones used in (Umeyama, 1988), which concludes our proof.
@@ 4.4 Alignment and Complexity analysis
To analyze the complexity of our approach we study the 3 main parts of T-GAE: a) The design of the input structural features, b) The message-passing GNN that produces node embeddings, and c) the linear assignment algorithm. The computation of our neighborhood-based structural features is expected to take $\mathcal{O}\left(\left|\mathcal{V}\right|\right)$ in real graphs, as proved in (Henderson et al., 2011). The computational and memory complexity of the message-passing GNN is $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c\right)$, and $\mathcal{O}\left(\left|\mathcal{V}\right|c\right)$, where $c$ is the width of the GNN. The computational complexity to align the nodes of the graph is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$ since we are using the suboptimal greedy Hungarian. If we want to optimally solve the linear assignment problem we need to use the Hungarian algorithm that has $\mathcal{O}\left(\left|\mathcal{V}\right|^{3}\right)$ complexity. If we want to process large graphs we can use efficient nearest neighbors algorithms with complexity $\mathcal{O}\left(\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ to perform soft linear assignment. Overall the complexity of T-GAE is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$, or $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c+\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ for large graphs.
@@
Arenas (Leskovec & Krevl, 2014) 1,133 5,451 5,451 Email Communication
Douban (Zhang & Tong, 2016) 3,906 7,215 7,215 Social Network
@@
Subraph Matching ACM-DBLP (Zhang & Tong, 2019) 9,872 9,916 39,561 44,808 6,352 Citation Network
Douban Online-Offline (Zhang & Tong, 2016) 3,906 1,118 1,632 3,022 1,118 Social Network
@@ 5 Experiments
In this section, we evaluate the performance of the proposed framework on both graph and sub-graph alignment with various benchmark networks. We compare against several baselines and assess the performance of the competing methods in terms of matching accuracy, hit-rate, and runtime.
@@ 5.1 Datasets and baselines
Table 2 provides a brief overview of the considered networks. Our comparisons are conducted with 3 categories of baseline methods: (a) GNN based methods: WAlign (Gao et al., 2021), GAE and VGAE (Kipf & Welling, 2016); (b) Graph/Node embedding techniques: NetSimile (Berlingerio et al., 2013), Spectral (Umeyama, 1988), DeepWalk (Perozzi et al., 2014), (Grover & Leskovec, 2016b), GraphWave (Donnat et al., 2018) and LINE (Tang et al., 2015). (c) Optimization based graph matching algorithms: S-GWL (Xu et al., 2019a), ConeAlign (Chen et al., 2020) and FINAL (Zhang & Tong, 2016). Note that LINE, VGAE, DeepWalk, and Node2Vec are omitted from some experiments since they show very poor performance. The reason behind that is that they are not permutation equivariant. GraphWave is also excluded from the sub-graph matching experiment, it could not identify correlated nodes in two different graphs. In the case of graphs without attributes FINAL is equivalent to the popular Isorank (Singh et al., 2008) algorithm. FINAL is omitted in sub-graph matching experiments due to weak performance.
@@ 5.2 Model Details
[…]
@@ 5.3.1 Transferability across networks
Table 3: Graph matching accuracy on 10 randomly perturbed samples under different levels of edge editing. The proposed T-GAE is trained on the clean Celegans, Arena, Douban, and Cora networks, and tested on noisy versions of them and the larger Dblp, and Coauthor CS. We test 3 different message-passing mechanisms for the layers of T-GAE as annotated in the table. Accuracy above 80% is highlighted in green, 40% to 80% accuracy is in yellow, and performance below 40% is in red.
@@
Spectral
Netsimile
GraphWave
FINAL
S-GWL
ConeAlign
WAlign
GAE
T-GAE
$0\%$ perturbation Celegans ${87.8\pm 1.5}$ ${72.7\pm 0.9}$ ${65.3\pm 1.7}$ ${92.2\pm 1.2}$
$\mathbf{93.0\pm 1.5}$
${66.6\pm 1.2}$
${88.4\pm 1.6}$
${80.2\pm 2.6}$
${88.9\pm 1.3}$
Arenas
$\mathbf{97.7\pm 0.4}$
${94.7\pm 0.3}$
${81.7\pm 0.7}$
${97.5\pm 0.3}$
${97.5\pm 0.3}$
${87.8\pm 0.6}$
${97.4\pm 0.5}$
${95.5\pm 0.4}$
${97.6\pm 0.4}$
Douban
${89.9\pm 0.4}$
${46.4\pm 0.4}$
${17.5\pm 0.2}$
${89.9\pm 0.3}$
$\mathbf{90.1\pm 0.3}$
${68.1\pm 0.4}$
${90.0\pm 0.4}$
${65.6\pm 0.4}$
${89.9\pm 0.4}$
Cora
${85.0\pm 0.4}$
${73.7\pm 0.4}$
${8.3\pm 0.4}$
$\mathbf{87.5\pm 0.7}$
${87.3\pm 0.7}$
${38.5\pm 0.7}$
${87.2\pm 0.4}$
${84.0\pm 0.8}$
${87.3\pm 0.4}$
Dblp
${84.5\pm 0.1}$
${63.7\pm 0.2}$
doesn’t scale
$\mathbf{85.6\pm 0.2}$
> 10 hours
${44.3\pm 0.6}$
${85.6\pm 0.2}$
${74.1\pm 0.3}$
$\mathbf{85.6\pm 0.2}$
Coauthor CS
${97.5\pm 0.1}$
${90.9\pm 0.1}$
doesn’t scale
$\mathbf{97.6\pm 0.1}$
> 10 hours
${75.8\pm 0.5}$
${97.5\pm 0.2}$
${71.1\pm 0.3}$
$\mathbf{97.6\pm 0.1}$
$1\%$ perturbation Celegans ${68.5\pm 16.1}$ ${66.3\pm 3.8}$ ${22.5\pm 22.4}$ ${33.2\pm 7.8}$
$\mathbf{92.4\pm 1.4}$
${60.9\pm 2.5}$
${80.7\pm 3.0}$
${13.6\pm 11.9}$
${84.1\pm 1.1}$
Arenas
${85.0\pm 10.0}$
${87.8\pm 1.0}$
${40.5\pm 23.8}$
${32.5\pm 5.9}$
${48.80\pm 48.7}$
${84.6\pm 1.0}$
${90.0\pm 3.1}$
${12.8\pm 12.5}$
$\mathbf{94.2\pm 0.8}$
Douban
${25.8\pm 27.2}$
${40.0\pm 1.2}$
${9.9\pm 5.9}$
${27.8\pm 5.7}$
${8.9\pm 26.7}$
${64.7\pm 0.4}$
${77.2\pm 4.8}$
${3.3\pm 2.9}$
$\mathbf{84.9\pm 0.6}$
Cora
${59.1\pm 9.3}$
${66.4\pm 1.6}$
${3.7\pm 2.9}$
${30.0\pm 3.3}$
${26.0\pm 39.6}$
${33.5\pm 1.6}$
${80.1\pm 1.2}$
${12.5\pm 5.1}$
$\mathbf{82.9\pm 0.5}$
Dblp
${55.6\pm 19.0}$
${55.1\pm 1.7}$
doesn’t scale
${15.2\pm 3.3}$
> 10 hours
${37.8\pm 1.1}$
${73.1\pm 1.6}$
${3.9\pm 0.9}$
$\mathbf{79.1\pm 0.4}$
Coauthor CS
${58.2\pm 22.1}$
${75.2\pm 2.2}$
doesn’t scale
${13.3\pm 5.0}$
> 10 hours
${68.5\pm 2.8}$
${75.2\pm 5.4}$
${3.0\pm 1.2}$
$\mathbf{86.5\pm 0.8}$
$5\%$ perturbation Celegans ${24.9\pm 15.9}$ ${41.1\pm 13.0}$ ${7.6\pm 9.2}$ ${10.4\pm 2.7}$
$\mathbf{81.1\pm 27.0}$
${50.5\pm 3.4}$
${42.4\pm 21.1}$
${3.9\pm 5.1}$
${48.3\pm 9.1}$
Arenas
${52.1\pm 16.5}$
${52.3\pm 5.3}$
${6.9\pm 7.2}$
${7.2\pm 2.6}$
${18.9\pm 37.7}$
$\mathbf{75.0\pm 2.7}$
${30.4\pm 17.5}$
${2.6\pm 2.6}$
${61.6\pm 8.0}$
Douban
${23.8\pm 20.6}$
${20.7\pm 4.6}$
${1.9\pm 2.8}$
${7.8\pm 3.0}$
${0.0\pm 0.0}$
${54.1\pm 1.2}$
${36.6\pm 13.4}$
${0.6\pm 0.6}$
$\mathbf{57.9\pm 6.1}$
Cora
${29.5\pm 0.8}$
${41.2\pm 3.3}$
${0.8\pm 0.3}$
${6.7\pm 2.8}$
${16.3\pm 32.6}$
${23.0\pm 2.0}$
${33.4\pm 7.3}$
${2.2\pm 0.9}$
$\mathbf{58.2\pm 2.0}$
Dblp
${28.0\pm 7.8}$
${19.5\pm 4.8}$
doesn’t scale
${2.7\pm 0.9}$
> 10 hours
${24.4\pm 2.9}$
${15.9\pm 8.3}$
${0.5\pm 0.1}$
$\mathbf{40.8\pm 2.1}$
Coauthor CS
${9.7\pm 5.0}$
${26.3\pm 6.0}$
doesn’t scale
${2.0\pm 0.4}$
> 10 hours
$\mathbf{51.4\pm 5.1}$
${11.3\pm 7.5}$
${0.1\pm 0.0}$
${26.9\pm 5.4}$
Our first observation is that for zero perturbation most algorithms are able to achieve a high level of matching accuracy. This is expected, since for zero perturbation the network alignment is equivalent to graph isomorphism. Furthermore, there is a clear benefit of processing the NetSimile embeddings with GNNs since they offer up to $22\%$ performance increase. When some perturbation is added, the conclusions are straightforward. Our proposed T-GAE markedly outperforms all the competing alternatives and shows the desired robustness to efficiently perform network alignment at $1\%$ perturbation level, and its performance is consistent across all datasets and perturbation levels.
Regarding the ability of T-GAE to perform large-scale network alignment the results are definitive. T-GAE enables low-complexity training with small graphs, and execution at larger settings by leveraging transfer learning. In particular, it is able to achieve very high levels of matching accuracy for both Dblp and Coauthor CS, for $p=0\%,~{}0.1\%$. To the best of our knowledge, this is the first attempt that performs exact alignment on a network at the order of $20$k nodes and $80$k edges.
Comparing T-GAE with vanilla GAE, we observe that GAE is not robust to noise or transferable. This highlights the benefit of T-GAE in handling the distribution shift brought by the structural dissimilarity between different graphs. We also compare runtime between T-GAE and all competing methods, and the results are presented in AppendixF.
5.3.2 Transferability across perturbation models

### v3

@@ 1 Introduction
Network alignment, also known as graph matching, is a classical problem in graph theory, that aims to find node correspondence across different graphs and is vital in a number of high-impact domains (Emmert-Streib et al., 2016). In social networks, for instance, network alignment has been used for user deanonymization (Nilizadeh et al., 2014) and analysis (Ogaard et al., 2013), while in bioinformatics it is a key tool to identify functionalities in protein complexes (Singh et al., 2008), or to identify gene–drug modules (Chen et al., 2018). Graph matching also finds application in computer vision (Conte et al., 2003), sociology (Racz and Sridhar, 2021), or politics (Li et al., 2022), to name a few. Graph matching can be cast as a quadratic assignment problem (QAP), which is in general NP-hard
Various approaches have been developed to tackle network alignment and can be divided into two main categories; i) optimization algorithms that attempt to approximate the QAP problem by relaxing the combinatorial constraints, ii) embedding methods that approach the problem by implicitly or explicitly generating powerful node embeddings that facilitate the alignment task. Optimization approaches, as (Anstreicher and Brixius, 2001; Vogelstein et al., 2015) employ quadratic programming relaxations, while (Klau, 2009) and (Peng et al., 2010) utilize semidefinite or Lagrangian-based relaxations respectively. Successive convex approximations were also proposed by (Konar and Sidiropoulos, 2020) to handle the QAP. Challenges associated with these methods include high computational cost, infeasible solutions, or nearly optimal initialization requirements. Embedding methods, on the other hand, overcome these challenges, but they usually produce inferior solutions, due to an inherent trade-off between embedding permutation-equivariance and the ability to capture the structural information of the graph. Typical embedding techniques include spectral and factorization methods (Umeyama, 1988; Feizi et al., 2019; Zhang and Tong, 2016; Kanatsoulis and Sidiropoulos, 2022), structural feature engineering methods (Berlingerio et al., 2013; Heimann et al., 2018), and random walk approaches (Perozzi et al., 2014; Grover and Leskovec, 2016a). Recently (Chen et al., 2020; Karakasis et al., 2021) have proposed joint node embedding and network alignment, to overcome these challenges, but these methods do not scale up as the size of the graph increases.
@@
Graph Neural Networks (GNNs) are powerful architectures that learn graph representations (embeddings). They have shown state-of-the-art performance in several tasks, including biology (Gainza et al., 2020; Strokach et al., 2020; Jiang et al., 2021), quantum chemistry (Gilmer et al., 2017), social networks and recommender systems (Ying et al., 2018; Wu et al., 2020; Liu et al., 2023; Yang et al., 2024). Recently, (Gao et al., 2021) proposed a GNN approach to match attributed graphs. The method used a joint embedding framework for pairs of graphs and achieved high levels of matching accuracy. However, this method does not scale to large graphs, since training graphs with large sizes is computationally prohibitive.
@@ 2.1 Network Alignment
As mentioned in the introduction, network alignment, is equivalent to the QAP, which has been proven to be NP-hard (Koopmans and Beckmann, 1957).
@@ 3 Graph Neural Networks (GNNs) Upper-Bounds Spectral Methods for Network Alignment
The above property is not satisfied by other spectral methods. GNNs are also stable (Gama et al., 2020; Parada-Mayorga et al., 2023; Ruiz et al., 2021), transferable (Ruiz et al., 2020), and have high expressive power (Xu et al., 2019b; Abboud et al., 2021; Kanatsoulis and Ribeiro, 2022).
@@ 3.1 GNNs and Network Alignment
The proof can be found in Appendix C. The assumption that the graph adjacencies have different eigenvalues is not restrictive. Real nonisomorphic graphs have different eigenvalues with very high probability (Haemers and Spence, 2004). Theorem 3.2 compares the network alignment power of a GNN with that of a spectral algorithm Umeyama (1988), that uses the absolute values of graph adjacency eigenvectors to match two different graphs. According to Theorem 3.2 there always exists a GNN that can perform at least as well as the spectral approach. The proof studies a GNN with white random input and measures the variance of the filter output. Then it shows that GNN layers are able to compute the absolute values of the graph adjacency eigenvectors when the adjacency has non-repeated eigenvalues. As a result there always exists a single layer GNN that outputs the same node features as the ones used in Umeyama (1988), which concludes our proof.
@@ 4.4 Alignment and Complexity analysis
To analyze the complexity of our approach we study the 3 main parts of T-GAE: a) The design of the input structural features, b) The message-passing GNN that produces node embeddings, and c) the linear assignment algorithm. The computation of our neighborhood-based structural features is expected to take $\mathcal{O}\left(\left|\mathcal{V}\right|\right)$ in real graphs, as proved in Henderson et al. (2011). The computational and memory complexity of the message-passing GNN is $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c\right)$, and $\mathcal{O}\left(\left|\mathcal{V}\right|c\right)$, where $c$ is the width of the GNN. The computational complexity to align the nodes of the graph is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$ since we are using the suboptimal greedy Hungarian. If we want to optimally solve the linear assignment problem we need to use the Hungarian algorithm that has $\mathcal{O}\left(\left|\mathcal{V}\right|^{3}\right)$ complexity. If we want to process large graphs we can use efficient nearest neighbors algorithms with complexity $\mathcal{O}\left(\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ to perform soft linear assignment. Overall the complexity of T-GAE is $\mathcal{O}\left(\left|\mathcal{V}\right|^{2}\right)$, or $\mathcal{O}\left(\left|\mathcal{V}\right|c^{2}+\left|\mathcal{E}\right|c+\left|\mathcal{V}\right|\text{log}\left(\left|\mathcal{V}\right|\right)\right)$ for large graphs.
@@
Arenas (Leskovec and Krevl, 2014) 1,133 5,451 5,451 Email Communication
Douban (Zhang and Tong, 2016) 3,906 7,215 7,215 Social Network
@@
Subraph Matching ACM-DBLP (Zhang and Tong, 2019) 9,872 9,916 39,561 44,808 6,352 Citation Network
Douban Online-Offline (Zhang and Tong, 2016) 3,906 1,118 1,632 3,022 1,118 Social Network
@@ 5 Experiments
The experiments included in this section are designed to answer the following research questions:(1)Can T-GAE generate competing graph matching accuracy while being efficient by utilizing transferability of GNN? (2)Is it empirically evident that node features processed by T-GAE is more robust to perturbations than existing GNN based framework? (3)Is T-GAE robust to different graph matching tasks and real-world noise distributions? (4)Is T-GAE generalizable to networks from various domains with different sizes?
To answer these questions, we conduct graph matching experiments by introducing different levels of edge editing under different perturbation models. The results are included in Section 5.3 and Appendix H. We also conducted sub-graph matching experiments to align two different real-world networks that are partially aligned, which are presented in Section 5.4. For each experiment, we include a collection of bench-marked datasets and competing baselines. We also compare the running time of different algorithms, the results are appended in Appendix F.
@@ 5.1 Datasets and baselines
Table 2 provides a brief overview of the considered networks. Our comparisons are conducted with 3 categories of baseline methods: (a) GNN based methods: WAlign (Gao et al., 2021), GAE and VGAE (Kipf and Welling, 2016); (b) Graph/Node embedding techniques: NetSimile (Berlingerio et al., 2013), Spectral (Umeyama, 1988), DeepWalk (Perozzi et al., 2014), (Grover and Leskovec, 2016b), GraphWave (Donnat et al., 2018) and LINE (Tang et al., 2015). (c) Optimization based graph matching algorithms: S-GWL (Xu et al., 2019a), ConeAlign (Chen et al., 2020) and FINAL (Zhang and Tong, 2016). Note that LINE, VGAE, DeepWalk, and Node2Vec are omitted from some experiments since they show very poor performance. The reason behind that is that they are not permutation equivariant. GraphWave is also excluded from the sub-graph matching experiment, it could not identify correlated nodes in two different graphs. In the case of graphs without attributes FINAL is equivalent to the popular Isorank (Singh et al., 2008) algorithm. FINAL is omitted in sub-graph matching experiments due to weak performance.
@@ 5.2 Model Details
As illustrated in Figure 1, the structure of our proposed encoder consists of two MLPs and a series of GNN layers. The node features are processed by a 2-layer MLP and passed to all the GNN layers. We add skip connections between this MLP layer and all the subsequent GNN layers. The outputs of all GNN layers are concatenated and passed to another 2-layer MLP, followed by a linear decoder to generate the reconstructed graph. The model is optimized end to end by equation 8. Note that one can choose different message passing functions as $f$ and $g$ in Equation (4). Our reported results are based on GIN(Xu et al., 2019b).
@@ 5.3 Graph Matching Experiments
Table 3: Graph matching accuracy on 10 randomly perturbed samples under different levels of edge editing on Uniform model. The proposed T-GAE is trained on the clean Celegans, Arena, Douban, and Cora networks, and tested on noisy versions of them and the larger Dblp, and Coauthor CS. Accuracy above 80% is highlighted in green, 60% to 80% accuracy is in yellow, and performance below 60% is in red.
@@
Spectral Netsimile GraphWave FINAL S-GWL ConeAlign WAlign GAE T-GAE
$0\%$ perturbation Celegans ${87.8\pm 1.5}$ ${72.7\pm 0.9}$ ${65.3\pm 1.7}$ ${92.2\pm 1.2}$ $\mathbf{93.0\pm 1.5}$ ${66.6\pm 1.2}$ ${88.4\pm 1.6}$ ${86.3\pm 1.3}$ ${91.0\pm 1.1}$
Arenas ${97.7\pm 0.4}$ ${94.7\pm 0.3}$ ${81.7\pm 0.7}$ ${97.5\pm 0.3}$ ${97.5\pm 0.3}$ ${87.8\pm 0.6}$ ${97.4\pm 0.5}$ ${97.6\pm 0.4}$ $\mathbf{97.8\pm 0.4}$
Douban ${89.9\pm 0.4}$ ${46.4\pm 0.4}$ ${17.5\pm 0.2}$ ${89.9\pm 0.3}$ $\mathbf{90.1\pm 0.3}$ ${68.1\pm 0.4}$ ${90.0\pm 0.4}$ ${89.5\pm 0.4}$ $\mathbf{90.1\pm 0.3}$
Cora ${85.0\pm 0.4}$ ${73.7\pm 0.4}$ ${8.3\pm 0.4}$ ${87.5\pm 0.7}$ ${87.3\pm 0.7}$ ${38.5\pm 0.7}$ ${87.2\pm 0.4}$ ${87.1\pm 0.8}$ $\mathbf{87.5\pm 0.4}$
Dblp ${84.5\pm 0.1}$ ${63.7\pm 0.2}$ doesn’t scale $\mathbf{85.6\pm 0.2}$ > 48 hours ${44.3\pm 0.6}$ ${85.6\pm 0.2}$ ${85.2\pm 0.3}$ $\mathbf{85.6\pm 0.2}$
Coauthor CS ${97.5\pm 0.1}$ ${90.9\pm 0.1}$ doesn’t scale $\mathbf{97.6\pm 0.1}$ > 48 hours ${75.8\pm 0.5}$ ${97.5\pm 0.2}$ ${97.6\pm 0.3}$ $\mathbf{97.6\pm 0.1}$
$1\%$ perturbation Celegans ${68.5\pm 16.1}$ ${66.3\pm 3.8}$ ${22.5\pm 22.4}$ ${33.2\pm 7.8}$ $\mathbf{87.1\pm 6.1}$ ${60.9\pm 2.5}$ ${80.7\pm 3.0}$ ${33.2\pm 8.4}$ ${86.5\pm 1.1}$
Arenas ${85.0\pm 10.0}$ ${87.8\pm 1.0}$ ${40.5\pm 23.8}$ ${32.5\pm 5.9}$ ${94.2\pm 0.7}$ ${84.6\pm 1.0}$ ${90.0\pm 3.1}$ ${30.1\pm 17.6}$ $\mathbf{96.0\pm 1.0}$
Douban ${25.8\pm 27.2}$ ${40.0\pm 1.2}$ ${9.9\pm 5.9}$ ${27.8\pm 5.7}$ ${72.1\pm 0.7}$ ${64.7\pm 0.4}$ ${77.2\pm 4.8}$ ${38.3\pm 16.4}$ $\mathbf{87.3\pm 0.4}$
Cora ${59.1\pm 9.3}$ ${66.4\pm 1.6}$ ${3.7\pm 2.9}$ ${30.0\pm 3.3}$ ${46.4\pm 6.9}$ ${33.5\pm 1.6}$ ${80.1\pm 1.2}$ ${57.9\pm 5.3}$ $\mathbf{85.1\pm 0.5}$
Dblp ${55.6\pm 19.0}$ ${55.1\pm 1.7}$ doesn’t scale ${15.2\pm 3.3}$ > 48 hours ${37.8\pm 1.1}$ ${73.1\pm 1.6}$ ${19.4\pm 0.6}$ $\mathbf{83.3\pm 0.4}$
Coauthor CS ${58.2\pm 22.1}$ ${75.2\pm 2.2}$ doesn’t scale ${13.3\pm 5.0}$ > 48 hours ${68.5\pm 2.8}$ ${75.2\pm 5.4}$ ${49.5\pm 7.8}$ $\mathbf{93.2\pm 0.8}$
$5\%$ perturbation Celegans ${24.9\pm 15.9}$ ${41.1\pm 13.0}$ ${7.6\pm 9.2}$ ${10.4\pm 2.7}$ ${68.3\pm 12.7}$ ${50.5\pm 3.4}$ ${42.4\pm 21.1}$ ${6.5\pm 2.4}$ $\mathbf{69.2\pm 2.1}$
Arenas ${52.1\pm 16.5}$ ${52.3\pm 5.3}$ ${6.9\pm 7.2}$ ${7.2\pm 2.6}$ ${68.0\pm 2.9}$ ${75.0\pm 2.7}$ ${30.4\pm 17.5}$ ${1.4\pm 1.4}$ $\mathbf{78.6\pm 2.5}$
Douban ${23.8\pm 20.6}$ ${20.7\pm 4.6}$ ${1.9\pm 2.8}$ ${7.8\pm 3.0}$ ${68.6\pm 0.8}$ ${54.1\pm 1.2}$ ${36.6\pm 13.4}$ ${0.6\pm 0.3}$ $\mathbf{70.2\pm 2.5}$
Cora ${29.5\pm 0.8}$ ${41.2\pm 3.3}$ ${0.8\pm 0.3}$ ${6.7\pm 2.8}$ ${39.9\pm 5.5}$ ${23.0\pm 2.0}$ ${33.4\pm 7.3}$ ${9.6\pm 2.7}$ $\mathbf{67.7\pm 1.3}$
Dblp ${28.0\pm 7.8}$ ${19.5\pm 4.8}$ doesn’t scale ${2.7\pm 0.9}$ > 48 hours ${24.4\pm 2.9}$ ${15.9\pm 8.3}$ ${1.4\pm 0.2}$ $\mathbf{60.8\pm 1.9}$
Coauthor CS ${9.7\pm 5.0}$ ${26.3\pm 6.0}$ doesn’t scale ${2.0\pm 0.4}$ > 48 hours ${51.4\pm 5.1}$ ${11.3\pm 7.5}$ ${0.6\pm 0.1}$ $\mathbf{66.0\pm 1.4}$
Our first observation is that for zero perturbation most algorithms are able to achieve a high level of matching accuracy. This is expected, since for zero perturbation the network alignment is equivalent to graph isomorphism. Furthermore, there is a clear benefit of processing the NetSimile embeddings with TGAE since it offers up to $43.7\%$ performance increase. When some perturbation is added, the conclusions are straightforward. Our proposed T-GAE markedly outperforms all the competing alternatives and shows the desired robustness to efficiently perform network alignment at $1\%$ and $5\%$ perturbation levels, and its performance is consistent across all datasets and perturbation levels.
Regarding the ability of T-GAE to perform large-scale network alignment the results are definitive. T-GAE enables low-complexity training with small graphs, and execution at larger settings by leveraging transfer learning. In particular, it is able to achieve very high levels of matching accuracy for both Dblp and Coauthor CS, for $p=0\%,~{}1\%$. It is also the only method that consistently achieves at least $60\%$ accuracy at $5\%$ perturbation. To the best of our knowledge, our experiments on DBLP (Pan et al., 2016) and Coauthor CS (Shchur et al., 2018) are the first attempts to perform exact alignment on networks at the order of $20$k nodes and $80$k edges.
Comparing T-GAE with vanilla GAE, we observe that GAE is not robust to noise. This highlights the benefit of T-GAE in handling the distribution shift brought by the structural dissimilarity between different graphs. We also compare runtime between T-GAE and all competing methods, and the results are presented in AppendixF.
5.4 Sub-graph Matching Experiments
Table 4: Subgraph matching performance comparison. The proposed T-GAE is trained on the two real-world graphs, and test to match the aligned portion of them.